OpenStudy (anonymous):

f(t)=t^2 simplify 4 [f(w-1)]+10

OpenStudy (anonymous):

f(t) = t^2 -> 4 [(w-1)^2]+10 = 4 [ w^2 -2w + 1 ] + 10 = 4w^2 - 8w + 14